f¹₃ (⨯)=cotx4.siny²x³
1
Expert's answer
2020-03-11T09:34:34-0400

"f_x=-\frac{4x^3}{\sin^2x^4}\sin(y^2x^3)+3x^2y^2\cos(y^2x^3)\cot x^4"

"f_y=2yx^3\cos(y^2x^3)\cot x^4"